The sum of three numbers is 94. The second number is 2 less than the first number, and the third number is 6 more than the first. Determine the numbers.
Found 2 solutions by solver91311, rfer:
Answer by solver91311(24713)



Let represent the first number.

Let represent the second number.

Let represent the third number.

Then:

1.

2.

3.

Substitute into 1:

1.

and solve for . The others can be easily calculated once you know the first one.

John


Answer by rfer(16322)

x+x-2+x+6=94
3x=90
x=30
x-2=28
x+6=36